# Scalable Decentralized Algorithms for Online Personalized Mean Estimation

### v3

@@ Appendix F Appendix - Proof of the results in Table 1
F.1 $n^{\star}_{\gamma}(x)$, sub-Gaussian setting
@@
From [Shalev-Shwartz and Ben-David, 2014, Lemma A.1] a sufficient condition for this inequality to hold is
@@
F.2 $n^{\star}_{\gamma}(x)$, bounded 4-th moment setting
@@
Appendix G Appendix - Additional Performance Evaluation of B-ColME and C-ColME
@@ G.1 Over a $G_{0}(N,r)$ varying $r$
G.2 B-ColME as a Function of the Depth of Information Kept
@@
G.3 C-ColME (Constant $\alpha$) as a Function of the Weight $\alpha$
@@
Appendix H Appendix - Additional Details on Decentralized Federated Learning (FL-DG)
@@
We focus on the case $C=2$ and use the MNIST dataset Deng [2012]. To obtain two different distributions from the MNIST dataset, we simply swap two labels (“3” and “5” as well as “1” and “7”). Each node has the task of recognizing handwritten digits using its local data and collaborating with neighboring nodes over $\mathcal{G}$, which is a complete graph in our scenario. We use a very simple feedforward neural network model for all nodes. It consists of the input layer, a hidden layer with $100$ nodes, and the output layer. We indicate the parameters of the NN as $\theta^{t}_{a}$, for agent $a$ at time $t$.
@@
Differently from the B-ColME and C-ColME mean estimation algorithms we have to modify the discovery phase to a large extent since the task of mean estimation and model training are structurally different. In Sattler et al. [2021] it was shown that it is possible to partition the agents in a federated learning framework by using the cosine similarity of the gradient updates (or the parameters updates) of the considered agents. Note that cosine similarity values close to $1$ indicate similar models/agents, while lower values indicate increasingly different agents. This can be intuitively understood by observing that two nodes with different data distributions are optimizing different loss functions, and, if we constrain the starting point of the optimization to be the same for both agents, we will observe an increase in the angle between the vectors corresponding to the gradient updates (see Figure 2 in Sattler et al. [2021] for an illustrative example). Subject to some regularity assumptions, it is indeed possible to use the cosine similarity of the parameter updates instead of gradients. Let us denote the updates as $\Delta\theta^{t}=\theta^{t+1}-\theta^{t}$.

### v4

@@ Appendix H Appendix H - Proof of the results in Table 2
H.1 $n^{\star}_{\gamma}(x)$, sub-Gaussian setting
@@
From (Shalev-Shwartz and Ben-David 2014, Lemma A.1) a sufficient condition for this inequality to hold is
@@
H.2 $n^{\star}_{\gamma}(x)$, Bounded 4-th Moment Setting
@@
Appendix I Appendix I - Additional Performance Evaluation of the B-ColME and C-ColME
@@ I.1 Over a $G_{0}(N,r)$ Varying $r$
I.2 B-ColME as a Function of the Depth of Information Kept
@@
I.3 C-ColME (Constant $\alpha$) as a Function of the Weight $\alpha$
@@
Appendix J Appendix J - Additional Details on the Decentralized Federated Learning Approach FL-DG
@@
We focus on the case $C=2$ and use the MNIST dataset (Deng 2012). To obtain two different distributions from the MNIST dataset, we simply swap two labels (“3” and “5” as well as “1” and “7”). Each node has the task of recognizing handwritten digits using its local data and collaborating with neighboring nodes over $\mathcal{G}$, which is a complete graph in our scenario. We use a very simple feedforward neural network model for all nodes. It consists of the input layer, a hidden layer with $100$ nodes, and the output layer. We indicate the parameters of the NN as $\theta^{t}_{a}$, for agent $a$ at time $t$.
@@
Differently from the B-ColME and C-ColME mean estimation algorithms we have to modify the discovery phase to a large extent since the task of mean estimation and model training are structurally different. In (Sattler, Müller, and Samek 2021) it was shown that it is possible to partition the agents in a federated learning framework by using the cosine similarity of the gradient updates (or the parameters updates) of the considered agents. Note that cosine similarity values close to $1$ indicate similar models/agents, while lower values indicate increasingly different agents. This can be intuitively understood by observing that two nodes with different data distributions are optimizing different loss functions, and, if we constrain the starting point of the optimization to be the same for both agents, we will observe an increase in the angle between the vectors corresponding to the gradient updates (see Figure 2 in (Sattler, Müller, and Samek 2021) for an illustrative example). Subject to some regularity assumptions, it is indeed possible to use the cosine similarity of the parameter updates instead of gradients. Let us denote the updates as $\Delta\theta^{t}=\theta^{t+1}-\theta^{t}$.